# AdaMesh: Personalized Facial Expressions and Head Poses for Adaptive Speech-Driven 3D Facial Animation

### v2

@@ 4 Experiments
Dataset. Due to the scarcity of 3D animation datasets with personalized talking styles, we utilize the 2D audio-visual talking face datasets. For expression adapter pre-train, we utilize Obama weekly footage [27] with a neutral talking style. For pose adapter pre-train, VoxCeleb2 [4] is utilized, which contains about 1M talking-head videos of 6112 speakers with diverse and natural head poses. In the adaptation stage, we introduce MEAD [33] for the expression adapter, since MEAD consists of videos featuring 60 actors with 7 different emotions. We utilize VoxCeleb2-Test dataset for the adaptation of the pose adapter, and the speakers are unseen in pre-train.
Dataset Processing. We resample 2D videos as 25fps and employ the 3D face reconstruction method [10, 11] to extract the identity, expression, and head pose. Given these three parameters, we can compute the 3D vertices of the per-frame meshes. For the audio track, we downsample the raw wave as 16kHz and adopt the base HuBERT model [13] to extract speech features. Since the frame rates of video and audio features are different, we interpolate the audio features and align them to the same length [5].
@@
FaceFormer 4.67 3.46 6.42 1.05 4.05$\pm$0.15 1.46$\pm$0.10 0.80%
Codetalker 4.43 3.30 6.48 1.05 4.10$\pm$0.14 1.42$\pm$0.11 0.80%
FACIAL 3.44 2.48 6.32 1.68 0.69 1.50 30.65 3.88$\pm$0.21 3.85$\pm$0.20 11.7%
GeneFace 3.54 2.47 6.56 1.38 0.70 1.49 34.13 4.01$\pm$0.18 3.79$\pm$0.17 12.3%
SadTalker 4.80 3.62 6.42 1.38 0.53 1.11 62.40 3.88$\pm$0.14 3.82$\pm$0.15 11.5%
Imitator 3.15 2.33 5.97 1.05 3.98$\pm$0.19 1.42$\pm$0.11 29.8%
Ground Truth 0 0 7.85 1.37 3.01 4.30$\pm$0.14 4.12$\pm$0.23 41.3%
AdaMesh (Ours) 2.91 2.25 7.10 0.90 0.87 1.57 13.89 4.12$\pm$0.19 4.22$\pm$0.11
Baseline Methods. We compare AdaMesh with several state-of-the-art methods. 1) FaceFormer [8], which first introduces autoregressive Transformer-based architecture for speech-driven 3D facial animation. 2) CodeTalker [34], which is the first attempt to model facial motion space with discrete primitives. 3)FACIAL [38], which focuses on predicting dynamic face attributes, including head poses. 4) GeneFace [35], which proposes a domain adaptation pipeline to bridge the domain gap between the large corpus and the target person video. It can also generate realistic head poses with a template pose sequence. 5) SadTalker [40], which learns the realistic 3D motion coefficient and generalize to various images. 6)Imitator [28], which learns identity-specific details from a short input video and produces facial expressions matching the personalized speaking style. FACIAL, GeneFace and SadTalker are photo-realistic talking face methods, we therefore reconstruct 3D meshes [10] from the generated videos or convert the BFM [7] to FLAME [16] format for comparison. More details can be found in the supplementary materials.
5 Results
5.1 Quantitative Evaluation
Evaluation Metric. To measure lip synchronization, we calculate the lip vertex error (LVE), which is widely used in previous methods [8, 28]. We follow [21] to calculate the emotional vertex error (EVE) to evaluate if the method can capture the personalized emotional talking style on MEAD dataset [33]. The diversity scores [25, 2] for expressions (Div-$\boldsymbol{\delta}$) and poses (Div-$\boldsymbol{\psi}$) are separately computed to evaluate if the methods generate vivid and diverse motions. The landmark standard deviation (LSD) [38, 34] measures the pose dynamics. The FID score [25, 41] estimates the head pose realism. To evaluate if the head poses are semantically aligned with the speech, we calculate the FSD score by computing the FID scores which correspond to the same HuBERT units and then averaging these scores  [13].
Evaluation Results. The results are shown in Tab.1. For facial expressions, it can be observed that our approach obtains lower LVE than other methods, which indicates AdaMesh has more accurate lip movements. FaceFormer and CodeTalker have poorer performance, since the LVE score is calculated on MEAD and Obama datasets, and these two methods have poor generalization ability to unseen speaker identity. AdaMesh also achieves the lowest EVE. It suggests that our approach can efficiently capture the facial style in the adaptation. A higher diversity score for expressions confirms that our approach generates more vivid facial dynamics. For head poses, FaceFormer, CodeTalker, and Imitator cannot generate head poses. GeneFace incorporates a real head pose sequence for presentation and  FACIAL learns over-smoothed poses. The FID scores show that our approach generates more realistic head poses, while LSD and diversity scores confirm our approach generates head poses with higher diversity. AdaMesh also gains the lowest FSD score, suggesting that the generated head poses are semantic aware.
Figure 4: Qualitative comparison with different methods. (a) shows lip movements on Obama dataset with neutral talking style. (b) is for observation of personalized facial expressions on the emotional MEAD dataset. (c) and (d) show head poses and corresponding landmark tracemaps on the VoxCeleb2-Test dataset. The first row displays the words or sentences that these frames are pronouncing.
@@ 5.2 Qualitative Results
Speech-Lip Synchronization. We illustrate three keyframes speaking the vowels and consonants. The visual results show that our approach generates more accurate lip movements than other methods. AdaMesh has more pronounced mouth open when speaking /ou/ and /ei/. FaceFormer and CodeTalker close their mouth tighter on /b/ than other methods, since they are directly trained on the 3D scanned vertices. Other methods are trained or fine-tuned on the reconstructed data from 2D video, which are less precise than the scanned data. AdaMesh also shows competitive performance on VOCASET dataset [5] and the details are described in the supplementary materials.
Facial Expression Richness. From Fig. 4(b), it can be observed that our approach preserves the expressive facial style and obviously gains richer facial details, especially around the cheek and lip regions. AdaMesh even gains more expressive details and accurate lip movements than the ground truth in the supplementary video. The MoLoRA parameters in AdaMesh act as a plugin of facial style for the pre-trained model, and the pre-trained model on the Obama dataset has guaranteed lip synchronization.
Head Pose Naturalness. As shown in Fig. 4(c), we illustrate frames in a sentence that contains evident emotion. A sentence that conveys negative semantics is accompanied by more evident head poses, which is consistent with the FSD result in the quantitative evaluation. Our approach produces head movements with a greater range of amplitudes, and the results are closer to the ground truth. Following previous works  [38, 40], we also plot the landmark tracemaps for a sentence that conveys negative semantics. The head poses with more degrees of freedom confirm the diversity generation of AdaMesh.
User Study. We conduct a user study to measure the perceptual quality of these methods. Fifteen users with professional English ability participate in this study. For lip synchronization and pose naturalness, we adopt mean-opinion-score (MOS) tests to ask users to score each video from one (worst) to five (best). Our approach has achieved the highest scores on these two aspects. For expression richness, we adopt A/B tests to ask users to select a better video between AdaMesh and another method. Our approach has significantly outperformed other methods, including the ground truth. This is attributed to the image jittering brought by [10, 16]. AdaMesh has filtered out these disturbances in the pre-training of the expression adapter. More user study details can be found in the supplementary materials.
@@ 5.3 Approach Analysis
Ablation Study for Expression Adapter. For expression adapter, we separately remove the following modules to verify their necessity. 1) As shown in Tab. 2, the removal of the style encoder leads to a drop of vertices errors and diversity metrics by a large margin. It suggests that the style encoder provides essential details which cannot be predicted from speech. 2) The absence of identity also causes a slight performance drop since the expression and identity parameters are not fully disentangled [16, 10]. The identity parameters also contain person-specific talking styles. 3) We fine-tune the full expression adapter instead of updating MoLoRA parameters. It can be observed that the performance significantly decreases since fine-tuning with few data damages the generalization ability of the pre-trained model.
Ablation Study for Pose Adapter. For pose adapter, we test three settings. 1) The retrieval of the pose style matrix is removed by randomly selecting a pose style embedding in the training data for inference. All metrics drop significantly due to the unawareness of the reference’s personalized talking style. 2) w/o HuBERT feature, where only the mel spectrogram is utilized for the speech representation. The LSD and FSD scores show that HuBERT features contain semantic information, which is vital to the diverse and meaningful head pose generation. 3) After removing VQ-VAE, the model directly predicts the continuous poses. The diversity scores drop to near zero. It suggests that VQ-VAE provides realistic quantized codes which is easier to predict. To further investigate the efficacy, we plot the yaw degree of head poses in Fig. 5(a). The pose curve of AdaMesh has the largest variation, which is closer to the ground truth. It confirms the results in quantitative and qualitative experiments. We also utilize t-SNE [31] to visualize the pose style matrix for the selected exciting and peaceful head poses. The visualization result shows that different styles are clustered into different groups. This proves that the semantic-aware pose style matrix contains meaningful pose style information.
Table 2: Ablation study for expression adapter.
@@
Table 3: Ablation study for pose adapter.
@@
Ours 0.90 0.87 1.57 13.89
w/o style matrix 1.01 0.69 0.69 19.88
w/o HuBERT feature 0.91 0.07 0.15 26.73
w/o VQ-VAE 0.91 0.05 0.02 22
Amount of Adaptation Data. We only analyze the impact of the adaptation data amounts on the expression adapter performance, since the pose adapter actually is a zero-shot model. As shown in Fig. 4(a), points within the same color family represent the same method, while the same marker corresponds to the computation of the same metric. Thus the points with the same marker can be compared. The figure has two y-axises, and the legend next to each y-axis indicates what metric this y-axis represents. We can conclude from this figure that, increasing the adaptation data amount can improve performance. AdaMesh achieves significantly better performance than Imitaor with the same amount of adaptation data.
Amount of Adaption Parameters. We estimate the impact of MoLoRA parameter amounts on the expression adapter performance. MoLoRA with different rank size combinations and the vanilla LoRA (the point in Fig. 4(b) with the maximum parameter amount) are compared. The point with the highest diversity score corresponds to the rank size combination of $[4,8,16,32]$. This setting also achieves the lowest LVE and EVE scores. Furthermore, when optimizing roughly the same number of parameters, our approach outperforms Imitator in all metrics.
Comparison with Other Topologies. As shown in Fig. 7, compared with other methods of different topologies that cannot be converted to FLAME format [16], AdaMesh still achieves better lip sync and richer facial expressions.
Figure 5: Expression adapter analysis. (a) Performance vs data amount. (b) Performance vs adaptation parameter amount.
Figure 6: Pose adapter analysis. (a) Yaw degree curves for head poses. (b) t-SNE visualization of the semantic-aware pose style matrices
Figure 7: Visual comparison with methods of different topologies.
@@ 6 Conclusion
7 Limitation
Limited by the adopted parametric face model [16], AdaMesh cannot freely and separately control the neck and head. This is inconsistent with the fact that human usually rotate their heads with relatively fixed shoulders. We will investigate modeling the neck dynamics in the future.

### v3

@@ 4 Experiments
Dataset. Due to the scarcity of 3D animation datasets with personalized talking styles, we utilize the 2D audio-visual talking face dataset. For expression adapter pre-train, we utilize Obama weekly footage (Suwajanakorn et al., 2017) with a neutral talking style. For pose adapter pre-train, VoxCeleb2 (Chung et al., 2018) is utilized, which contains about 1M talking-head videos of different celebrities with diverse and natural head poses. In the adaptation stage, we introduce MEAD (Wang et al., 2020) for the expression adapter, since MEAD consists of videos featuring 60 actors with 7 different emotions. We introduce VoxCeleb2 test dataset for the adaptation of the pose adapter.
Dataset Processing. We resample the 2D videos as 25fps and employ the 3D face reconstruction method DECA (Feng et al., 2021) to extract the identity, expression, and head pose. Given these three parameters, we can compute the 3D vertices of the per-frame meshes. For the audio track, we downsample the raw wave as 16000Hz and adopt the base HuBERT model (Hsu et al., 2021) to extract speech features. Since the frame rates of video and audio features are different, we interpolate the audio features and align them to the same length (Cudeiro et al., 2019).
Baseline Methods. We compare AdaMesh with several state-of-the-art methods. 1) FaceFormer (Fan et al., 2022), which first introduces autoregressive Transformer-based architecture for speech-driven 3D facial animation. 2) CodeTalker (Xing et al., 2023), which is the first attempt to model facial motion space with discrete primitives. 3) GeneFace (Ye et al., 2023), which proposes a domain adaptation pipeline to bridge the domain gap between the large corpus and the target person video. It can also generate realistic head poses with a template pose sequence. 4)Imitator (Thambiraja et al., 2023), which learns identity-specific details from a short input video and produces facial expressions matching the identity-specific speaking style. 5)FACIAL (Zhang et al., 2021), which focuses on predicting dynamic face attributes, including head poses. GeneFace and FACIAL are photo-realistic talking face methods, we therefore reconstruct 3D meshes Feng et al. (2021) from the generated videos for comparison.
5 Results
5.1 Quantitative Evaluation
Evaluation Metric. To measure lip synchronization, we calculate the lip vertex error (LVE), which is widely used in previous methods (Fan et al., 2022; Thambiraja et al., 2023). Since MEAD dataset contains significant emotion, we follow Peng et al. (2023) to calculate the emotional vertex error (EVE) to evaluate if the method can capture the personalized emotional talking style. The diversity scores (Siyao et al., 2022; Chen et al., 2023) for expressions (Div-$\bm{\delta}$) and poses (Div-$\bm{\psi}$) are separately computed to evaluate if the methods generate vivid and diverse motions. The landmark standard deviation (LSD) (Zhang et al., 2021; Xing et al., 2023) is used to measure the pose dynamics. The FID score (Siyao et al., 2022; Zhu et al., 2023) for motion sequence is also employed to estimate the pose realism.
Evaluation Results. The results are shown in Tab.1. For facial expressions, it can be observed that our method obtains lower LVE than other methods, which indicates AdaMesh has more accurate lip movements. FaceFormer and CodeTalker have poorer performance, since the LVE score is calculated on MEAD and Obama datasets, and these two methods have poor generalizability to unseen speaker identity. AdaMesh also achieves the lowest EVE. It suggests that our approach can efficiently capture the facial style in the adaptation. A higher diversity score for expressions confirms that our approach generates more vivid facial dynamics. For head poses, FaceFormer, CodeTalker, and Imitator cannot generate head poses. GeneFace incorporates a real head pose sequence for presentation and FACIAL learns over-smoothed poses. The FID scores show that our approach generates more realistic head poses, while LSD and diversity scores confirm our approach generates head poses with higher diversity.
@@ 5.1 Quantitative Evaluation
FaceFormer 4.67 3.46 6.42 1.05 0.0 0.0 4.05$\pm$0.15 1.46$\pm$0.10 0.80%
Codetalker 4.43 3.30 6.48 1.05 0.0 0.0 4.10$\pm$0.14 1.42$\pm$0.11 0.80%
FACIAL 3.44 2.48 6.32 1.68 0.69 1.50 3.88$\pm$0.21 3.85$\pm$0.20 11.7%
GeneFace 3.54 2.47 6.56 1.38 0.70 1.49 4.01$\pm$0.18 3.79$\pm$0.17 12.3%
Imitator 3.15 2.33 5.97 1.05 0.0 0.0 3.82$\pm$0.18 1.42$\pm$0.11 29.8%
Ground Truth 7.85 1.37 3.01 4.30$\pm$0.14 4.12$\pm$0.23 41.3%
AdaMesh (Ours) 2.91 2.25 7.10 0.90 0.87 1.57 4.12$\pm$0.19 4.22$\pm$0.11
@@ 5.2 Qualitative Results
Speech-Lip Synchronization. To evaluate the lip synchronization performance, we illustrate three keyframes speaking the vowels and consonants. The visual results show that our approach generates more accurate lip movements than other methods. AdaMesh has more pronounced mouth open when speaking /ou/ and /ei/. FaceFormer and CodeTalker close their mouth tighter on /b/ than other methods, even the ground truth, since they are directly trained on the 3D scanned vertices. Other methods are trained or fine-tuned on the reconstructed data from 2D video, which are less precise than the scanned data.
Facial Expression Richness. To evaluate if the methods can capture the facial talking style from the reference video. We compare GeneFace, Imitator, and our AdaMesh on MEAD dataset. It can be observed that our approach preserves the expressive facial style and obviously gains richer facial details, especially around the cheek and lip regions. AdaMesh even gains more expressive details and accurate lip movements than the ground truth. The MoLoRA parameters in AdaMesh act as a plugin of facial style for the pre-trained model, and the pre-trained model on a larger scale data has guaranteed lip synchronization.
Head Pose Naturalness. We adapt our pose adapter on the VoxCeleb2-Test dataset. As shown in Fig. 3(c), we illustrate frames in a sentence that contains evident emotion. A sentence that conveys strong emotions is accompanied by more evident head poses. Our approach produces head movements with a greater range of amplitudes, and the results are closer to the ground truth. Following FACIAL (Zhang et al., 2021), we also plot the landmark tracemaps for a sentence that conveys negative semantics. The head poses with more degrees of freedom confirm the diversity of AdaMesh.
User Study. We conduct a user study to measure the perceptual quality of these methods. Fifteen users with professional English ability participate in this study. For lip synchronization and pose naturalness, we adopt mean-opinion-score (MOS) tests to ask users to score each video from one (worst) to five (best). Our approach has achieved the highest score on these two aspects. For expression richness, we adopt A/B test to ask users to select a better video between AdaMesh and another method. Our approach has significantly outperformed other methods, including the ground truth. This is attributed to the image jittering brought by Feng et al. (2021); Li et al. (2017). AdaMesh has filtered out these disturbances in the pre-training of the expression adapter.
Figure 3: Qualitative comparison with different methods. (a) shows lip movements on Obama dataset with neutral talking style. (b) is for observation of facial expressions on the emotional MEAD dataset. (c) and (d) show head poses and corresponding landmark tracemaps on the VoxCeleb2-Test dataset. The first row displays the words or sentences that these frames are pronouncing.
@@ 5.3 Approach Analysis
Ablation Study for Expression Adapter. For expression adapter, we separately remove the style encoder, identity encoder, and MoLoRA to verify the necessity of these modules. 1) As shown in Tab. 3, the removal of the style encoder leads to a drop of vertices errors and diversity metrics by a large margin. It suggests that the style encoder provides essential details which cannot be predicted from speech. 2) The absence of identity also causes a slight performance drop since the expression and identity parameters are not fully disentangled (Li et al., 2017; Feng et al., 2021). The identity parameters also contain person-specific talking style. 3) We fine-tune the full expression adapter instead of updating MoLoRA parameters. It can be observed that the performance significantly decreases since fine-tuning with few data damages the generalization ability of the pre-trained model.
Ablation Study for Pose Adapter. For pose adapter, we test three settings. 1) The retrieval of the pose style matrix is removed by randomly selecting a pose style embedding in the training data for inference. All metrics drop significantly due to the unawareness of the reference’s personalized talking style. 2) w/o HuBERT feature, where only the mel spectrogram is utilized for the speech representation. The LSD and diversity scores show that HuBERT features contain semantic information, which is vital to the diverse head pose generation. 3) After removing VQ-VAE, the model directly predicts the continuous pose values. The LSD and diversity scores drop to near zero. It suggests that VQ-VAE provides realistic quantized codes which is easier to predict. To further investigate the efficacy, we plot the yaw degree of head poses in Fig. 4(c). The pose curve of AdaMesh has the largest variation, which is closer to the ground truth. It confirms the results in quantitative and qualitative experiments. We also utilize t-SNE (Van der Maaten & Hinton, 2008) to visualize the pose style matrix for the selected exciting and peaceful head poses. The visualization result shows that different styles are clustered into different groups. This proves that the semantic-aware pose style matrix contains meaningful pose style information.
Amount of Adaptation Data. We only analyze the impact of the amounts of adaptation data on the performance of the expression adapter, since the pose adapter actually is a zero-shot model with the retrieval strategy. Fig. 4(a) demonstrates the performance vs. adaptation data amount for AdaMesh. Points within the same color family represent the same method, while the same marker corresponds to the computation of the same metric. Thus the points with the same marker can be compared. There are two y-axis in the figure, and the legend next to each y-axis indicates what metric this y-axis represents. We can conclude from this figure that, increasing the amount of adaptation data can obviously improve the performance. AdaMesh achieves significantly better performance than Imitaor with the same amount of adaptation data.
Amount of adaption Parameters. To better estimate the impact of the amounts of MoLoRA parameters on the performance of the expression adapter, we evaluate the performance of MoLoRA with different rank size combinations and the vanilla LoRA (the point in Fig. 4(b) with the maximum number of parameters). The point with the highest diversity score corresponds to the rank sizes of $[4,8,16,32]$. This setting also achieves the lowest LVE and EVE scores. Furthermore, when optimizing roughly the same number of parameters, our approach outperforms Imitator in all metrics.
Table 2: Ablation for expression adapter.
@@
Table 3: Ablation for pose adapter.
@@
Ours 0.90 0.87 1.57
w/o style matrix 1.01 0.69 0.69
w/o HuBERT feature 0.91 0.07 0.15
w/o VQ-VAE 0.91 0.05 0.02
Figure 4: Approach analysis. (a) performance vs. amount of adaptation data. (b) performance vs. amount of trainable parameters. (c) yaw degree curves for head poses. (d) t-SNE visualization of the semantic-aware pose style matrices.
@@ 6 Conclusion
Appendix A Appendix
This supplemental material contains three subsections: the details of the expression adapter and the pose adapter in AdaMesh, implementations of baseline methods, and additional information about the user study.
A.1 Details of expression adapter and pose adapter
The architecture details of AdaMesh are presented in Tab. 4. The overall architecture of the expression adapter is illustrated in Fig. 1 of the main paper. For the identity encoder, $T$ frames identity parameters are fed into two convolutional layers and average to a vector with a dimension of 256. The speech sequence and expression sequence are separately processed by the audio encoder and style encoder. These two encoders are three conformer layers to output features with $T\times 256$ dimension. The encoded features are concatenated with the identity latent feature and sent into the expression decoder. We also adopt three Conformer layers for the decoder to output the predicted expression sequence. This is a simple architecture, but with robust and efficient performance, which is mainly attributed to the identity encoder and the style encoder providing the residual style information.
For the pose adapter, following Siyao et al. (2022), VQ-VAE downsample and upsample the pose sequence with a symmetrical network structure of 1D convolutional layers. We also adopt the cross-conditional causal attention mechanism other than the vanilla attention in Transformer, since the head poses of different degrees are entangled with each other. We present the loss function of VQ-VAE as below:
$\mathcal{L}_{VQ}=\mathcal{L}_{rec}(\hat{\bm{\psi}},\bm{\psi})+\left\|sg[\hat{z}]-z_{q}\right\|+\beta\left\|z-sg[z_{q}]\right\|$ (4)
where $\hat{\bm{\psi}}$ and $\bm{\psi}$ denotes output euler angle sequence and input ground truth, respectively. And the specific formula for $\mathcal{L}_{\text{rec}}$ can be expressed as:
$\mathcal{L}_{\text{rec }}(\hat{\bm{\psi}},\bm{\psi})=\left\|\hat{\bm{\psi}}-\bm{\psi}\right\|_{1}+\alpha_{1}\left\|\hat{\bm{\psi}}^{\prime}-\bm{\psi}^{\prime}\right\|_{1}+\alpha_{2}\left\|\hat{\bm{\psi}}^{\prime\prime}-\bm{\psi}^{\prime\prime}\right\|_{1}$ (5)
where $\bm{\psi}^{\prime}$ and $\bm{\psi}^{\prime\prime}$ represent the 1st-order (velocity) and 2nd-order (acceleration) partial derivatives of euler angle sequence $\bm{\psi}$ on time. In addition, $\alpha_{1}$, $\alpha_{2}$ and $\beta$ denote trade-off coefficients while $sg[\cdot]$ denotes ”stop gradient”.
Table 4: Architecture parameter list for AdaMesh.
Expression Adapter Identity Encoder Conv1d Kernel 5
Conv1d Channel Size 256
Identity Embedding Width 256
Audio Encoder Conformer Layer 3
Conformer Layer Head 4
Conformer Layer Embedding Width 256
Audio Embedding Width 256
Style Encoder Conformer Layer 3
Conformer Layer Head 4
Conformer Layer Embedding Width 256
Audio Embedding Width 256
Expression Decoder Conformer Layer 3
Conformer Layer Head 4
Conformer Layer Embedding Width 256
Audio Embedding Width 256
Conv1d Kernel 5
Conv1d Channel Size 256
Pose Adaptor VQ-VAE Encoder Layers 3
Decoder Layers 3
Encoder/Decoder Conv1d Kernel 4
Encoder/Decoder Conv1d Channel Size 256
Codebook Size 1024
Codebook Embedding Width 256
Pose GPT Transformer Layers 6
Transformer Layer Head 12
Transformer Layer Embedding Width 768
Style Embedding Width 50
A.2 Implementations of baseline methods
As mentioned in the main paper, we compare AdaMesh with state-of-the-art methods. FaceFormer (Fan et al., 2022) and CodeTalker (Xing et al., 2023) have released the pre-trained model, thus we test inference with the available codes. To calculate metrics with the same identity topology, we adopt Li et al. (2017) to recover the shape parameter from 2D video are convert it to the format that FaceFormer and CodeTalker require for inference. For GeneFace (Ye et al., 2023), we extract the required data to train the image render and fine-tune the post-net. For FACIAL (Zhang et al., 2021), since it is a person-specific method which requires a large amount of data for training, we directly utilize the pre-trained checkpoint of Obama for inference. For Imitator (Thambiraja et al., 2023), it is the most-related and latest work. We implement it to the best of our understanding. As soon as Imitator releases the public codes, we will replace the visualized images in our paper with the original ones.
A.3 User Study
The designed user study of the mean-of-score tests for lip synchronization and head pose naturalness and the A/B tests for the richness of facial expressions are separately demonstrated in Fig. 5 and Fig. 6. Prior to commencing the tests, we have thoroughly communicated the experimental guidelines and important information that users need to be aware of. For the lip synchronization and head pose naturalness tests, we sample 105 videos for seven compared methods. For the A/B tests, 60 pairs of compared videos are created.
Figure 5: Designed mean-of-score test interface for lip synchronization and head pose naturalness.
Figure 6: Designed A/B test interface for the richness of facial expressions.